Which of the following is a reflection of f(x) = -x + 1 about the x-axis?
a. r(x) = x + 1
b. r(x) = x - 1
c. r(x) = -(x - 1)
d. None of the above

Answer by stanbon(75887)

Which of the following is a reflection of f(x) = -x + 1 about the x-axis?
a. r(x) = x + 1
b. r(x) = x - 1
c. r(x) = -(x - 1)
d. None of the above
---------------------------
You are told "the y values are -x+1".
You want the negative of those y values;
you want -(-x+1) = x -1, which is answer "b"
================
Cheers,
Stan H.